# PLaD: Preference-based Large Language Model Distillation with Pseudo-Preference Pairs

### v1

@@ 2 Related Work
Learning from Preference Data. The pivotal Reinforcement Learning from Human Feedback (RLHF) framework (Christiano et al., 2017) first uses preference data to fit a reward model and then fine-tunes the LM to maximize the given reward using reinforcement learning algorithms. In practice, however, using RL to directly fine-tune LMs are challenging, incurring significant computational costs and requiring extensive hyperparameter tuning. To mitigate this issue, DPO (Rafailov et al., 2023) proposes to directly train the policy LM using the pairwise logistic loss without explicitly fitting a reward model. SLiC (Zhao et al., 2023) and RRHF (Yuan et al., 2023) adopt the pairwise hinge loss to train the policy LM. RSO (Liu et al., 2023) further introduces the statistical rejection sampling method to improve DPO by addressing the distribution drift issue. More recently, LiPO Liu et al. (2024) leverages a ranked list of responses (either rated by humans for ranked by models Qin et al. (2023)) for LLM alignment. All these studies require either human annotations or external reward models to obtain the preference data and focus on aligning a single model with human preference. In this work, we focus on distilling a teacher model into a student model with self-supervised preference pairs.
@@ 4.3 Distillation with Preference Pairs
The margin calibration loss is introduced to further refine the student’s output by considering a scoring function $s$, which provides an additional quality measure for the generated sequences:
$\begin{split}L^{\text{cal}}_{\text{margin}}&=\max(0,\beta(s(y;\hat{y}_{+};x)-s(y;\hat{y}_{-};y;x))\\ &-\log P_{\theta}(\hat{y}_{+}|x)+\log P_{\theta}(\hat{y}_{-}|x))\end{split}$ (5)
@@ 5.2 Main Results
Table 1: Main results with the LLaMA-2 model and GPT-Neo model families. For the LLaMA-2 group, the teacher model is LLaMA-2-13B, and the student model is LLaMA-2-7B. For the GPT-Neo group, the teacher model is GPT-Neo-2.7B, and the student model is GPT-Neo-1.3B. WR stands for win rate, and RM refers to a task-specific reward model. We also present the teacher-student win rate (T-S WR) on the distillation set evaluated by the reward model for reference.

### v2

@@ 2 Related Work
Learning from Preference Data. The pivotal Reinforcement Learning from Human Feedback (RLHF) framework (Christiano et al., 2017) first uses preference data to fit a reward model and then fine-tunes the LM to maximize the given reward using reinforcement learning algorithms. In practice, however, using RL to directly fine-tune LMs are challenging, incurring significant computational costs and requiring extensive hyperparameter tuning. To mitigate this issue, DPO (Rafailov et al., 2023) proposes to directly train the policy LM using the pairwise logistic loss without explicitly fitting a reward model. SLiC (Zhao et al., 2023) and RRHF (Yuan et al., 2023) adopt the pairwise hinge loss to train the policy LM. RSO (Liu et al., 2023) further introduces the statistical rejection sampling method to improve DPO by addressing the distribution drift issue. More recently, LiPO Liu et al. (2024) leverages a ranked list of responses (either rated by humans for ranked by models Qin et al. (2023)) for LLM alignment. All these studies require either human annotations or external reward models to obtain the preference data and focus on aligning a single model with human preference. Additionally, some works have attempted to explain the distillation mechanisms from multiple perspectivesLopez-Paz et al. (2015); Lopes et al. (2017); Zhang et al. (2020); Menon et al. (2021); Zhang et al. (2022), but none have yet bridged the preference learning and knowledge distillation, while it is a natural approach given the capacity gap between teacher and student models. In this work, we focus on distilling a teacher model into a student model with self-supervised preference pairs.
@@ 4.3 Distillation with Preference Pairs
The margin calibration loss is introduced to refine the student’s output by considering a scoring function $s$, which provides an additional quality measure for the generated sequences:
$\begin{split}L^{\text{cal}}_{\text{margin}}&=\max(0,\beta(s(y;\hat{y}_{+};x)-s(y;\hat{y}_{-};x))\\ &-\log P_{\theta}(\hat{y}_{+}|x)+\log P_{\theta}(\hat{y}_{-}|x))\end{split}$ (5)
@@ 5.2 Main Results
Table 1: Main results with the LLaMA-2 and GPT-Neo models. For the LLaMA-2 group, the teacher model is LLaMA-2-13B, and the student model is LLaMA-2-7B. For the GPT-Neo group, the teacher model is GPT-Neo-2.7B, and the student model is GPT-Neo-1.3B. WR stands for win rate, and RM refers to a task-specific reward model. The teacher-student win rate (T-S WR) on the distillation set evaluated by a reward model is provided for reference.
@@ A.6 Hyper-parameters
A.7 Score Function
TL;DR $\mathcal{L}_{\text{cal}}^{\text{rank}}$ TL;DR $\mathcal{L}_{\text{cal}}^{\text{margin}}$ HH $\mathcal{L}_{\text{cal}}^{\text{rank}}$ HH $\mathcal{L}_{\text{cal}}^{\text{margin}}$
LLaMA-2 40.46 40.21 27.65 27.74
GPT-Neo 14.70 14.76 10.38 9.94
Table 7: Ranking calibration loss v.s. margin calibration loss.
The score function in Eq. 5 is defined as
$s(\hat{y},y;x)=\sum_{n}F_{n}\left(e(\hat{y},y),e(\hat{y},x)\right),$ (6)
where $F_{n}=2P_{n}R_{n}/(P_{n}+R_{n})$. The definitions of $P_{n}$, $R_{n}$, $F_{n}$ can be found in (Zhang et al., 2019). When $n>1$, we have
$R_{n}=\frac{1}{|e|}\sum_{i+n\in e}\max_{j+n\in e}e_{i}^{T}e_{j+n},$ (7)
and
$P_{n}=\frac{1}{|e|}\sum_{j:n\in e}e_{i}^{T}e_{j:n+n}.$ (8)
This score function measures the similarity between the positive (negative) text and the reference text with negligible computational overhead, because it uses the student model to obtain the representation instead of an external model. By integrating it into Eq. 5, we aim to scale the margin $\beta$ with the score difference $s(\hat{y},y^{\prime};x)-s(\hat{y},y^{\prime\prime};x)$. This modulation increases the margin when the positive sequence closely resembles the reference text. A larger margin means that the positive sequences should be preferred more via a higher likelihood in the generation process. This dynamic adjustment of the margin essentially encourages a clearer distinction between positive and negative sequence pairs during training.
For the empirical validation, we choose between the ranking calibration loss and the margin calibration loss based on their performance on the validation set. We report the numbers of each loss in the main experiments setting here.